# Identifying Super Spreaders in Multilayer Networks

### v1

@@ 5.2 Result analysis
We begin the discussion of results by analysing the average performance of the evaluated methods, as illustrated by the cumulative relative score curves in Fig. 4 and the corresponding metrics in Tab. 3. ts-net achieved the best results in $7$ out of $8$ scenarios, significantly outperforming all other methods, particularly on real-world networks under the $T_{val}$ metric, which captures the ability to identify the most influential spreader for the entire graph. The only case in which ts-net was not ranked first was for real networks under the $S_{val}$ metric, where deg-cd held a slight advantage (by $0.004$). Among the baselines, only deg-cd and deg-c delivered comparable results, while all other methods exhibited substantially lower performance. These findings are further supported by Fig. 4, which also reveals a marked drop in performance across most methods beyond the top-ranked actors, indicating room for further improvement and investigation.
@@
Table 3: Average performance of the evaluated methods across two network splits. Predictions with the highest quality (i.e. closest to $1$) for each criterion are highlighted in green, with darker shades indicating better ranks.
@@
Table 4: Performance of the evaluated methods on the real-world networks. Predictions with the highest quality (i.e. closest to $1$) for each criterion are highlighted in green, with darker shades indicating better ranks. $S_{auc}$ covers mostly with $S_{val}$ and is hidden for conciseness.
@@
Despite this setup, ts-net was the only method to consistently achieve scores no lower than $0.790$ across all metrics and datasets. Moreover, it achieved the best performance in $11$ out of the $21$ evaluation cases, outperforming all other methods; the second-best method, deg-cd, ranked first in $9$ cases. ts-net obtained the highest $T_{val}$ score on five out of eight networks and consistently ranked among the top three in the remaining metrics. Its performance was especially pronounced on timik, demonstrating that predicting local rankings to improve memory efficiency does not compromise predictive quality. Notable performance deviations were observed only on the arxiv and eu-trans networks, both of which have considerably more layers ($13$ and $37$, respectively) than the maximum of $5$ layers seen in the training data.
Among the other methods, deg-cd and deg-c yielded stable and competitive performance, especially on smaller networks, where they frequently ranked in the top three. However, they underperformed on sparser or more irregular graphs such as timik. The results of nghb-s and nghb-sd were mixed: while effective in $T_{val}$ on arxiv, ckmp, and lazega, their scores in $S_{val}$ and $F_{auc}$ were inconsistent. The random baseline predictably placed last in most cases. mn2v-km performed reasonably on ckmp and timik, but fell behind top methods overall. Finally, DeepIM performed poorly with exception for $T_{val}$ on the small l2-course network and failed to run on the largest graph (timik) due to memory issues.
@@ 6 Conclusions
This study addresses the problem of identifying super-spreaders in multilayer networks by proposing a novel, generalisable solution. We introduced TopSpreadersNetwork, an inductive graph neural network capable of processing multilayer networks of varying sizes and complexities. As part of this work, we also constructed a dedicated dataset: TopSpreadersDataset, which captures actor-level diffusion characteristics under the Multilayer Independent Cascade Model. Finally, we proposed and validated alternative techniques for data transformation and embedding fusion that improve the robustness of ranking-learning task from multirelational data.
Despite these contributions, some limitations remain. While multilayer networks constitute a powerful and flexible framework, they are still a subset of the broader class of heterogeneous graphs, which are not directly addressed in this work. Moreover, although TopSpreadersNetwork demonstrates superior performance in our evaluations, relatively simple heuristics such as Degree Centrality continue to yield strong results. The main advantage of our approach lies in its interpretability: rather than producing scores in a black-box fashion, it first predicts a vector of four interpretable features, which can then be used to construct tailored rankings. This offers flexibility for use cases where, for instance, actors causing the highest peak of infections are more desirable than those triggering longer diffusions, something heuristic or end-to-end solutions do not easily allow.
Several promising directions exist for future research. First, the dataset may be extended in two independent ways. One option is to include diffusion results for small seed sets, generated using heuristics such as the greedy algorithm, thereby extending the learning task from scoring individual actors to evaluating entire seed sets. Another avenue is to reforge the dataset into a general benchmark for super-spreader identification, encompassing a broader range of diffusion models and network topologies.
@@
This research was partially supported by the National Science Centre, Poland, grant [no. 2022/45/B/ST6/04145], Higher Education, “International Projects Co-Funded” program, and the EU under the Horizon Europe, grant [no. 101086321]. However, views and opinions expressed here, are those of the authors only and do not necessarily reflect those of the funding agencies.
Appendix A Appendix: Model Selection and Ablation Study
@@ A.3 Data transformations
Appendix B Appendix: Evaluation on other metrics
In addition to evaluating the proposed model against competitive baselines using the relative cumulative score (Def. 5), we also employed metrics commonly used in the domain of ranking-learning. Instead of assessing prediction quality based on the scores assigned to top-$k$ ranked actors, these metrics evaluate only the alignment between the indices in $\mathbf{\hat{R}}$ and $\mathbf{R}$. Specifically, we report:
@@ Appendix B Appendix: Evaluation on other metrics
We report these metrics, similarly to the main metrics used in the study (Tab. 2), at three levels of the ranking: at its top position ($T$), at the cut-off separating super-spreaders ($S$), as illustrated in Fig. 1, and over the full ranking ($F$). For each metric, we report the average score computed across all prefixes of the ranking, that is, over all top-$k$ sets for $k=1,2,\dots,\{T,S,F\}$.
Table 12: Average precision of the evaluated methods across two network splits. Predictions with the highest quality ($\sim 1$) are highlighted in green, with darker shades indicating better ranks.
@@
Table 13: Average Jaccard index of the evaluated methods across two network splits. Predictions with the highest quality ($\sim 1$) are highlighted in green, with darker shades indicating better ranks.
@@
Appendix C Appendix: Training Details
All ts-net variants were trained using the AdamW optimiser with a OneCycleLR learning rate scheduler. The implementation was developed in Python 3.12 using PyTorch 2.3.1. Training was limited to a maximum of $50$ epochs, with early stopping applied and a patience threshold of $30$ epochs. NeighborLoader with a batch size of $128$, using an induced $N$-hop neighbourhood sampling strategy aligned with the number of convolutional layers was used. Specifically, up to $32$ neighbours were sampled at the first hop, with the number halved at each subsequent hop. The maximum learning rate was set to $0.003$. All experiments involving ts-net were conducted on a machine equipped with an Intel(R) Xeon(R) Gold 6330 CPU @ 2.00GHz, 64 GB RAM, and an NVIDIA Titan RTX GPU. Hyperparameters for baseline machine learning methods were configured as described in their respective original papers and trained on CPU due to high memory requirements.

### v2

@@ 5.2 Result analysis
We begin the discussion of results by analysing the average performance of the evaluated methods, as illustrated by the cumulative relative score curves in Fig. 4 and the corresponding metrics in Tab. 3. ts-net achieved the best results in 777 out of 888 scenarios, significantly outperforming all other methods, particularly on real-world networks under the $T_{val}$ metric, which captures the ability to identify the most influential spreader for the entire graph. The only case in which ts-net was not ranked first was for real networks under the $S_{val}$ metric, where deg-cd held a slight advantage (by 0.0040.0040.004). Among the baselines, only deg-cd and deg-c delivered comparable results, while all other methods exhibited substantially lower performance. These findings are further supported by Fig. 4, which also reveals a marked drop in performance across most methods beyond the top-ranked actors, indicating room for further improvement and investigation.
@@
Table 3: Average performance of the evaluated methods across two network splits. Predictions with the highest quality (i.e., closest to 111) for each criterion are highlighted in green, with darker shades indicating better ranks.
@@
Table 4: Performance of the evaluated methods on the real-world networks. Predictions with the highest quality (i.e., closest to 111) for each criterion are highlighted in green, with darker shades indicating better ranks. $S_{auc}$ covers mostly with $S_{val}$ and is hidden for conciseness.
@@
Despite this setup, ts-net was the only method to consistently achieve scores no lower than 0.7900.7900.790 across all metrics and datasets. Moreover, it attained the best performance in 111111 out of the 212121 evaluation cases, outperforming all other methods; the second-best method, deg-cd, was ranked first in 999 cases. ts-net obtained the highest $T_{val}$ score on five out of eight networks and consistently ranked among the top three in the remaining metrics. Its performance was especially pronounced on timik, demonstrating that predicting local rankings to improve memory efficiency does not compromise predictive quality. Notable performance deviations were observed on the arxiv and eu-trans networks, both having considerably more layers (131313 and 373737, respectively) than the maximum of 555 in the training data. The performance drop on out-of-distribution networks may stem from limitations of the aggregation component. While scalability w.r.t. the number of actors has been achieved using NeighborLoader, the number of relationships requires further investigation. The observed effect similar to that of oversquashing (too much information from many layers is compressed into one vector) could be mitigated using techniques analogous to those employed for handling large neighbourhoods; i.e., by processing the layers “chunk by chunk”, rather than all at once. Naturally, enriching the dataset with networks containing more layers may also help.
Among the other methods, deg-cd and deg-c yielded stable and competitive performance, especially on smaller networks, where they frequently ranked in the top three. However, they underperformed on more irregular graphs such as timik. The results of nghb-s and nghb-sd were mixed: while effective in $T_{val}$ on arxiv, ckmp, and lazega, their scores in $S_{val}$ and $F_{auc}$ were inconsistent. The random, predictably, placed last in most cases. mn2v-km performed reasonably on ckmp and timik, but fell behind top methods overall. Finally, DeepIM performed poorly, except for $T_{val}$ on the smallest l2-course, and failed to run on the largest graph (timik) due to memory issues.
@@ 6 Conclusions
This study addresses the problem of identifying super-spreaders in multilayer networks by proposing a novel solution — TopSpreadersNetwork — an inductive graph neural network capable of processing multilayer networks of varying sizes and complexities. To this end, we also constructed the TopSpreadersDataset, which captures actor-level diffusion characteristics under the Multilayer Independent Cascade Model. Finally, we proposed and validated alternative techniques for data transformation and embedding fusion that improve the robustness of ranking-based learning from multirelational data.
Nonetheless, some limitations remain. While multilayer networks constitute a powerful and flexible framework, they are still a subset of the broader class of heterogeneous graphs, which are not directly addressed in this work. Moreover, although TopSpreadersNetwork demonstrates superior performance, relatively simple heuristics such as Degree Centrality continue to yield strong results (for a comparison of time performance see App. C). The main advantage of our approach lies in its interpretability: rather than producing scores in a black-box fashion, it first predicts a vector of four interpretable features, which can then be used to construct tailored rankings. This offers flexibility for use cases where, e.g, actors causing the highest peak of infections are more desirable than those triggering longer diffusion, something other solutions do not easily allow.
Several directions exist for future research. One interesting option is to include diffusion results for small yet effective seed sets, e.g., generated using the greedy algorithm, thereby extending the learning task from scoring individual actors to evaluating entire sets. Another promising avenue is to reforge the dataset into a general benchmark for super-spreader identification, encompassing a broader range of diffusion models and network topologies.
@@
This research was partially supported by the National Science Centre, Poland, grant no. 2022/45/B/ST6/04145, the Polish Ministry of Science and Higher Education programme: International Projects Co-Funded, and the EU under the Horizon Europe, grant no. 101086321. Views and opinions expressed here are those of the authors only and do not necessarily reflect those of the funding agencies.
Appendix A Model selection and ablation study
@@ A.3 Data transformations
Appendix B Evaluation on other metrics
[…]
@@ Appendix B Evaluation on other metrics
[…]
Table 12: Average precision of the evaluated methods across two network splits. Predictions with the highest quality (1$\sim$ 1∼ 1) are highlighted in green, with darker shades indicating better ranks.
@@
Table 13: Average Jaccard index of the evaluated methods across two network splits. Predictions with the highest quality (1$\sim$ 1∼ 1) are highlighted in green, with darker shades indicating better ranks.
@@
Appendix C Scalability of the model
To assess the scalability of the proposed model in terms of computational efficiency, two aspects were analysed: GPU memory consumption and processing time as a function of the network size. The evaluation was based on the test set, as it was also processed by the competing methods considered in this study.
Figure 5: Inference time of ts-net compared to that of deg-c.
First, Fig. 5 presents processing times depending on the average number of edges per network layer for ts-net and, for reference, the light yet effective deg-c. The results for the proposed model indicate that processing time increases almost linearly, from near-zero values for small networks to approximately 161616 seconds for networks containing over 60,00060,00060 , 000 edges per layer. In the case of deg-c, the scaling is also linear but with a gentler slope. Linear regression trends confirm these observations: ts-net exhibits a slope of 0.00020.00020.0002 with a strong fit ($R^{2}=0.853$), while deg-c shows a slightly lower slope of 0.000090.000090.00009 with $R^{2}=0.860$. It is worth noting that comparisons with deg-c on larger graphs would not be representative due to presumably lower-level optimisation mechanisms in the degree function from the NetworkX library, which underpins this heuristic.
Regarding the second aspect, thanks to the use of NeighborLoader, the GPU memory usage of ts-net during inference remains constant at around 15.1615.1615.16 GB across the entire test set. This enables effective scaling and ensures applicability to real-world scenarios involving continuously growing networks.
Appendix D Training details
All ts-net variants were trained using the AdamW optimiser with a OneCycleLR learning rate scheduler. The implementation was developed in Python 3.12 using PyTorch 2.3.1. Training was limited to a maximum of 505050 epochs, with early stopping applied and a patience threshold of 303030 epochs. NeighborLoader with a batch size of 128128128, using an induced $N$-hop neighbourhood sampling strategy aligned with the number of convolutional layers was used. Specifically, up to 323232 neighbours were sampled at the first hop, with the number halved at each subsequent hop. The maximum learning rate was set to 0.0030.0030.003. All experiments involving ts-net were conducted on a machine equipped with an Intel(R) Xeon(R) Gold 6330 CPU @ 2.00GHz, 64 GB RAM, and an NVIDIA Titan RTX GPU. Hyperparameters for baseline machine learning methods were configured as described in their respective original papers and trained on CPU due to high memory requirements.